Alex cuts a 36 inch piece of wire into 2 pieces. Each piece will be bent into a square. The area of the larger square will be four times the area of the smaller square. What will be the area of the smaller square?
Answer by rwm(914)

a/4*a/4=4*(b/4*b/4)
a^2/16=b^2/4
4a^2=16b^2
a^2=4b^2
a=2b
a=-2b
we reject any solutions that are negative.
a+b=36
2b+b=36
3b=36
b=12
a=36-b
a=24
a=24 b=12
each side of the 24 inch wire will be 6 inches
6*6=36 in^2
each side of the 12 in wire will be 3 inches
3*3=9 in^2
check
36 is 4 times 9
ok